OpenStudy (anonymous):

Given that log10 (x-y+1)=0 and 1+log10 (xy)=0, show that x=y=1/sqrt of 10.

ganeshie8 (ganeshie8):

change the equations into exponent form and solve

ganeshie8 (ganeshie8):

use below : \(\large \log_a b = c\) means \(\large b = a^c\)

OpenStudy (anonymous):

do i need put the 1/sqrt of 10 into each of the equation?

ganeshie8 (ganeshie8):

nope, start by changing the given two equations to exponent form

ganeshie8 (ganeshie8):

\(\large \log_{10} (x-y+1)=0\) means \(\large x-y + 1 = 10^0\) \(\large x-y + 1 = 1\) \(\large x-y = 0 ~~~~~~~~~~ \color{Red}{(1)}\)

ganeshie8 (ganeshie8):

see if you can convert the second log equation ?

ganeshie8 (ganeshie8):

Alright, here it is : \( \large 1+\log_{10} (xy)=0\) \( \large \log_{10} (xy)=-1\) \( \large xy =10^{-1}\) \( \large xy =\dfrac{1}{10}~~~~~~\color{red}{(2)}\)

ganeshie8 (ganeshie8):

solve both the converted equations \(\color{red}{(1)}\) and \(\color{red}{(2)}\)

OpenStudy (anonymous):

oh , i got it thanks , but what does you mean solve? just substitute 1/sqrt 10 with x and y?

ganeshie8 (ganeshie8):

no, you need to solve both equations and show that x = 1/sqrt(10) and y = 1/sqrt(10)

ganeshie8 (ganeshie8):

\(\large x - y = 0\) \(\large xy = \frac{1}{10}\)

ganeshie8 (ganeshie8):

first equation gives you \(x = y\) plug that in second equation and solve \(y\)

OpenStudy (anonymous):

ok i know it now , thanks for your guide

ganeshie8 (ganeshie8):

good to hear :) can you show me the remaining work ?

OpenStudy (anonymous):

from first equation x=y , y^2= 1/10 y=1/sqrt (10) is it correct?

ganeshie8 (ganeshie8):

Excellent !!!

OpenStudy (anonymous):

really thank you so much

ganeshie8 (ganeshie8):

np, you're welcome :)